For the following reaction, Kc=255 at 1000 K.


CO(g)+Cl2(g) ⇌ COCl2(g)


A reaction mixture initially contains a CO concentration of 0.1550 M and a Cl2 concentration of 0.176 M at 1000 K.

You may want to reference (Pages 656-664) Section 15.8 while completing this problem.



Part A

What is the equilibrium concentration of CO at 1000 K ?

Part B

What is the equilibrium concentration of Cl2 at 1000 K ?


Part C

What is the equilibrium concentration of COCl2 at 1000 K ?
Answer #1

molar conc. at equilibrium. E Huw - x tx X . CO3+ .cg = Cocco, Ke = 255 at 1000k: For this reaction kc = [cochl = 255, wherecon = 0.155 - 3C = 0.155 M - 0.1398 m. eqm = 0.0152M 1921egm = 0.176-32 = 0.176 -0.1398) M = 0.0362 M. och eam = x = 0.1398 M